MC CH6 #29 B & C

MC CH6 #29 B & C

MC CH6 #29 B & C
Answer #1

B) For the reaction

CH4 --> C(g) + 4H

The enthalpy of reaction = Sum of enthalpy of products- sum of enthalpy of reactants

Enthalpy of reaction = (enthalpy of C(g) + 4X enthalpy of H) - (Enthalpy of CH4)

Let us put the values from table given

Etnhalpy of reaction = (718.4 + 4 X 217.94) - (-74.8) = 1664.96 KJ / mole


C) For reverse reaction

Enthalpy of reaction = (Enthalpy of CH4) - (enthalpy of C(g) + 4X enthalpy of H)

Enthalpy of reaction = -74.8 - ( 718.4 + 4X 217.94) = -1664.96 KJ / mole